OpenStudy (anonymous):

A piece of aluminum occupies a volume of 14.0 milliliters and weighs 32.6 grams. What is its density of the aluminum rounded to the nearest hundredth? Only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

@camerondoherty

OpenStudy (inkyvoyd):

what's the formula for density?

OpenStudy (inkyvoyd):

D=M/V.

OpenStudy (inkyvoyd):

M=32.6 V=14 do algebra

OpenStudy (inkyvoyd):

units will be in g/ml

OpenStudy (inkyvoyd):

also no need to tag 10 people at a time it's super annoying don't ever do that, asking in chat is good enough

OpenStudy (anonymous):

mass/volume = density

OpenStudy (hanner_b_nanner):

D=32.6/14... what does that give you?

OpenStudy (inkyvoyd):

^

OpenStudy (anonymous):

2.3 @hanner_B_nanner

OpenStudy (hanner_b_nanner):

Ok, that's your answer! :D

OpenStudy (anonymous):

yall n!ggas crae crae thing i can do this sh!t i hate math im salty when it comes to math lmfao

OpenStudy (anonymous):

can u help me with another one? @hanner_B_nanner

OpenStudy (hanner_b_nanner):

I can try :)

OpenStudy (anonymous):

A large emerald with a mass of 982.7 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

yeah me 2 lol @Tazmaniadevil

OpenStudy (anonymous):

lmfao

OpenStudy (anonymous):

So do you think you could help me or no? @hanner_B_nanner

OpenStudy (hanner_b_nanner):

ignore that...v=m/d v=982.7/2.76

OpenStudy (anonymous):

356 or 356.1 ?

OpenStudy (anonymous):

@hanner_B_nanner

OpenStudy (hanner_b_nanner):

356.05 because it says the hundredth place.

OpenStudy (anonymous):

ok thanks so much @hanner_B_nanner

OpenStudy (hanner_b_nanner):

anytime :D

OpenStudy (anonymous):

but wait it says round to the nearest hundreth

OpenStudy (hanner_b_nanner):

after the decimal it goes 10's place, hundreds place....

OpenStudy (anonymous):

so the answer would still be 356.05 ?

OpenStudy (hanner_b_nanner):

Do you have like multiple choice answers?

OpenStudy (anonymous):

oh okay lol thanks!

OpenStudy (anonymous):

and no i dont

OpenStudy (hanner_b_nanner):

ohhh, that stinks :( lol

OpenStudy (anonymous):

yes it does! :(